OpenStudy (anonymous):

Use the distributive property to simplify the expression below: The quantity of 3 times x minus 21 all divided by 3

OpenStudy (anonymous):

\[\left( 3x -21 \right) \over 3\]?

OpenStudy (anonymous):

Is this the expression you mean?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok... write \(\Large \frac{3x-21}{3} \) as \(\Large \frac{1}{3}(3x-21) \) so you can distribute the 1/3 to the parenthesis...

OpenStudy (anonymous):

but i think it would be better to simplify the expression by factoring the numerator as 3(x-7) and cancel common factors from numerator/denominator...